x+2y=3
find th eslope and y intercept for each line the has a slpoe and y intercept

Answer by checkley71(8403)

X+2Y=3 OR 2Y=-X+3 OR Y=-X/2+3/2 THE SLOPE IS (-1/2) AND THE Y INTERCEPT IS (3/2)